the floor of a shed has an area of 90 square feet. the floor is in the shape of a rectangle whose length is 3 feet less than twice the width. find the length and the width of the floor of the shed. use the formula area= length x width
Answer by Boreal(15235)

l*w=90
x=width
2x-3=length
x(2x-3)=90 units are ft*ft=sq ft.
2x^2-3x-90=0, moving the 90 to the left side.
(2x-15)(x+6)=0
2x-15=0
x=(15/2), or 7.5 ft
The other foot is negative and may be ignored.
Width is 7.5 feet
Length is 12 feet.
Area is 90 sq ft.